OpenStudy (anonymous):

If two sides of a triangle are 9 cm and 15 cm in length, which COULD be the measure of the third side?

OpenStudy (anonymous):

if we divide both sides by 3 we have 3 and 5 this is a triangle 3-?-5

OpenStudy (anonymous):

ok how do i find the missing number

OpenStudy (anonymous):

 pythagoras

OpenStudy (anonymous):

yes ,, he's right .. use the pythagorean theorem to solve that triangle ..

OpenStudy (anonymous):

so 5^2 + 3^2= c^2

OpenStudy (anonymous):

no .. it's not the equation ..

OpenStudy (anonymous):

23,24,25,26

OpenStudy (anonymous):

15^2=9^2+x^2 225-81=x^2 144=x^2 x=12

OpenStudy (anonymous):

in normal cases you cud use the rate of a 3-4-5 triangle we noticed that we have 3-?-5 so its 3-4-5 hancere verse the calculation made b4 by multiplying by 3 4(3)=12

OpenStudy (anonymous):

o the answer is ?

OpenStudy (anonymous):

12

OpenStudy (anonymous):

thns